OpenStudy (anonymous):

To the nearest tenth of a percent, what is 8/13written as a percent?

hartnn (hartnn):

to get anything in % multiply it by 100 x = x*100%

hartnn (hartnn):

so, 8/13 = 800/13 = .... (use calculator)

OpenStudy (anonymous):

8/13 = .62 = 62%